OpenStudy (awy):

Car A is traveling at 18 m/s and car B at 25 m/s. Car A is 300m behind car B when the driver of Car A accelerates his car with an acceleration of 1.80m/s^2. How long does it take car A to overtake car B?

OpenStudy (anonymous):

Use the equation \[s= \frac{1}{2}at^2 +ut\] Where u=initial velocity

OpenStudy (anonymous):

s= distance; a= acceleration , and t =time

OpenStudy (anonymous):

Rearrange the equation, then solve.

OpenStudy (awy):

ok, is s like the intial distance?

OpenStudy (anonymous):

s = total distance

OpenStudy (anonymous):

In this case, the total distance that car A will travel is 300m, hence s=300m

OpenStudy (awy):

so, i got 300t + 0.9t^2

OpenStudy (awy):

wait nvm, its actually 0.9t^2 + u^2 - 300 ?

OpenStudy (anonymous):

Close ... check the u^2 ..

OpenStudy (awy):

ohh. 0.9^2 + 18^t - 300

OpenStudy (awy):

0.9t^2 rather

OpenStudy (anonymous):

Yup. Good.

OpenStudy (anonymous):

Make it equal to 0, then solve using the quadratic formula.

OpenStudy (awy):

\[\frac{ -18\pm \sqrt{32.9} }{ 1.8 }\]

OpenStudy (anonymous):

Check the number which is being square-rooted.

OpenStudy (awy):

oh there shouldnt be a sq root over 32.9

OpenStudy (anonymous):

It should be \[\sqrt{b^2-4ac}\]

OpenStudy (awy):

\[\sqrt{18^{2}-4(0.9)(-300)}\]

OpenStudy (anonymous):

Yep. Now put it into the quadratic formula

OpenStudy (anonymous):

And remember, your answer should be positive, because you're trying to find the time, and time is always positive.

OpenStudy (awy):

ok, so it should look like this? \[\frac{ -18\pm37.5 }{ 1.8 }\]

OpenStudy (awy):

and the positive answer is 10.8 secs?

OpenStudy (anonymous):

Correct. =)

OpenStudy (awy):

Thank you very much :P

OpenStudy (anonymous):

You're welcome!